WATER BROADENING OF OXYGEN

\underline{BRIAN J. DROUIN}, VIVIENNE PAYNE, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109; ELI MLAWER, Atmospheric and Environmental Research, 131 Hartwell Avenue, Lexington, MA 02421.

A need for precise air-mass retrievals utilizing the near-infrared O_{2} A-band has motivated measurements of the waterbroadening in oxygen. Experimental challenges have resulted in very little water broadened oxygen data, especially in the near-infrared where pressure broadened linewidth must compete with the relatively large thermal linewidth. Existing water broadening data^{a} for the O_{2} A-band is of insufficient precision for application to the atmospheric data. Because of the nature of scattering processes, it is believed that broadening parameters for O_{2} from one spectral region may be transferable to other spectral regions - so we investigated the O_{2} 60 GHz magnetic dipole Q branch which is also used prominently in remote sensing. Atmospheric retrievals of air-mass and temperature that use the 60 GHz magnetic dipole Q branch incorporate a water-broadening parameter that is scaled to self-broadened values, but there is only high temperature data that directly supports this hypothesis.^{b} We present precise O_{2}-H_{2}O broadening measurements for the magnetic dipole Q-branch and the pure-rotational band, measured at room temperature with a Zeeman-modulated absorption cell and a frequency-multiplier spectrometer. Here we will describe the apparatus and the measurement analysis. Inter-comparisons of these and other O_{2} broadening data sets confirm the expectation of only minor band-to-band scaling of pressure broadening. The measurement provides a basis for fundamental parameterization of retrieval codes for the long-wavelength atmospheric measurements. Finally, we encourage the application of these measurements for retrievals of air-mass via remote sensing of the oxygen A-band. a E.M. Vess et al. J. Phys. Chem. A 116, 4069-4073 (2012). b G. Fanjoux et al. J. Chem. Phys. 101(2) 1061-1071 (1994).